# Dynamical cluster-based optimization of tensor network algorithms for quantum circuit simulations

### v1

@@ III Entanglement clustering in finite-fidelity simulations
The core idea of this work is to improve the algorithms for simulating quantum circuits described in the previous section using cluster-based protocols. By leveraging the properties of a given circuit’s structure, we can “cluster” qubits based on regions of the circuit with a large number of entangling gates. This physics-informed clustering approach allows TEBD to adaptively divide the circuit into multi-layered clusters, employing optimized tensor network contractions [21] to contract each cluster exactly. In DMRG, a similar approach is applied through an adaptive clustering algorithm, which exploits hypergraph partitioning [35] to automatically form clusters of qubits, based on the distribution of entangling operations within a specified segment of the circuit of up to $L_{\mathrm{max}}$ gate layers. This qubit grouping is used to generate a grouped MPS, whose sites are variationally optimized in each iteration. Our implementation of the DMRG algorithm is adaptive, as the structure of the grouped MPS dynamically changes at each iteration to match the entangling gate distribution in the local circuit region.
@@
Figure 6: Schematic of the iteration process of the cluster-TEBD on an $8$-qubit, $7$-layer quantum circuit, with $Q_{\mathrm{max}}=5$. (a) Up to layer $2$, the circuit forms $4$ clusters of entanglement, with two qubits each. At layer $3$, the operation $U_{(6,7),3}$ groups clusters $3$ and $4$ in a new cluster $3$ with $4$ qubits. As the size of this cluster does not violate the condition on $Q_{\mathrm{max}}$, we move on to the next layer. (b) At layer $4$, operations $U_{(1,2),4}$ and $U_{(5,6),4}$ are performed on pairs of qubits that are already in the same cluster. Consequently, the clusters remain the same and we move on to the next layer. (c) At layer $5$, the operation $U_{(4,5),5}$ groups clusters $2$ and $3$ into a new cluster $2$ with 6 qubits. As the size of this cluster violates the condition on $Q_{\mathrm{max}}$, layer $4$ is the chosen layer up to which the circuit can be contracted exactly. (d) Schematic of the division of clusters up to layer $4$, contraction and decomposition - with truncation, if necessary. The obtained MPS is used as the initial state of the remaining evolution.
@@ III.1 Cluster-TEBD
The cluster-TEBD algorithm modifies the TEBD algorithm described in Sec. II.3 to allow the exact contraction of multiple gate layers of a quantum circuit. The idea is to divide the network in clusters of entanglement, based on the structure of the circuit to simulate. These clusters are independent from each other, meaning they can be interpreted as intermediate-to-large-sized tensor networks that can be exactly contracted in an efficient way.
@@
The idea of gate clustering has already been successfully applied in the context of exact simulations of quantum circuits [32, 33, 34], and our purpose is to exploit it in finite-fidelity tensor network algorithms.
@@
In the context of exact tensor network contractions, a similar approach is the “slicing” technique [21, 40, 41, 42, 43, 44, 45, 34, 46, 47, 48], where a network is divided into independent subnetworks by choosing a subset of indices to fix. As a consequence, the contraction of all subnetworks has to be repeated multiple times, for every possible value of the fixed indices. The main difference in our approach is that no indices within the network are fixed, eliminating the need to repeat the contraction of the generated subnetworks. Instead, the circuit partitioning is only based on the entanglement distribution and on the available memory, making sure that, after all subnetworks have been contracted, the resulting tensor sizes are kept manageable for the SVD compression step.
@@
A single iteration of cluster-TEBD is performed as follows. We define an array of length $N-1$, representing bonds between neighboring qubits, with all entries initially set to False. Each entry $i$ in the array corresponds to the bond between qubits $i$ and $i+1$, and its value indicates whether an entangling operation has been applied between them. The algorithm iterates over all gates in the circuit, starting from the layer closest to the initial state, to track the distribution of entangling operations. When encountering an $n$-qubit gate, with $n>1$, acting on qubits $i,\dots,i+n-1$, the corresponding entries $i,\dots,i+n-2$ of the bond array are updated to True. This signifies that an entangling operation has been applied between these qubits. At the end of each layer, the algorithm identifies entanglement clusters by finding consecutive entries in the bond array that are set to True. In the context of the cluster-TEBD algorithm, an entanglement cluster can be interpreted as a portion of the circuit, up to the current layer $L_{\text{current}}$, where qubits frequently interact with each other but never interact with the rest of the system. Since distinct clusters of entanglement can be treated independently, the corresponding circuit portion can be contracted as a separate circuit. The size of the tensor resulting from the contraction of a cluster depends on two quantities:
@@
Knowing that all the physical indices of a quantum circuit have dimension $2$, the size of tensor $\Psi^{[k]}_{L_{\text{current}}}$ that exactly encodes the entanglement cluster $k$ up to layer $L_{\text{current}}$ can be computed as:
@@
The quantum circuit can then be contracted exactly up to the last layer $L$ that satisfies both Eqs. (7) and (8). Once layer $L$ has been identified, the algorithm proceeds to exactly contract all entanglement clusters, dividing each portion of the circuit from the initial state up to $L$ in independent tensor networks. The exact contraction of each cluster $k$ exploits the greedy algorithm to find the corresponding contraction sequences [36, 21]. The resulting tensors $\Psi_{L}^{[k]}~{}\in~{}\mathbb{C}^{\chi_{q_{1}^{[k]}-1}\times q_{1}^{[k]}\times\dots\times q_{n}^{[k]}\times\chi_{q_{n}^{[k]}}}$ are finally decomposed using successive SVDs - as shown back in Fig. 2.
@@
Figure 7: Grouping algorithm for DMRG on an $8$-qubit quantum circuit, setting $L_{\mathrm{max}}=7$. (a) Generation of the virtual MPS from the quantum circuit by counting the number of two-qubit gates per bond and estimating the bond dimensions. (b) Recursive bipartitioning routine for the virtual MPS, setting $Q_{\mathrm{max}}=4$. The generated partitioning is finally used to group the initial state accordingly, creating the grouped MPS for the DMRG algorithm.
III.2 Enhancing DMRG
The idea of clustering together qubits based on the distribution of entanglement in a quantum circuit can also be used to improve the DMRG algorithm from Sec. II.3. As opposed to regular, periodically invariant circuit structures, for example a brick-wall structure with alternating two-qubit gates, generic circuits with irregular structures display an uneven distribution of entanglement throughout the evolution. Depending on the time step, distinct subsets of qubits may experience numerous entangling operations, while other separate subsets may undergo very few or even no operations. Our goal is to develop entanglement-based protocols that enhance the applicability of DMRG for simulating generic quantum circuits.
In Sec. II.3, we introduced the concept of a grouping scheme, designed to group together highly-entangled sites in the simulated quantum circuit. This idea of forming a grouped MPS is conceptually similar to creating clusters of entanglement, with one key difference: in the original DMRG algorithm, the grouping was predetermined prior to the algorithm’s execution. While this method works well for circuits with regular structures, it becomes less effective for deep circuits with irregularly arranged entangling gates, where it is not straightforward to predict how the bond dimensions will grow in different portions of the circuit. Therefore, we introduce a clustering approach to adaptively choose the site grouping and the needed bond dimensions, according to the structure of the simulated circuit.
An iteration of the adaptive grouping algorithm works as follows. Given a generic $N$-qubit quantum circuit, we count the number of entangling operations $E_{L_{\mathrm{max}}}(b)$ for each bond $b$, up to the chosen number of contracted layers per iteration $L_{\mathrm{max}}$. Then, we create a virtual MPS with $N$ sites, where each bond dimension is defined as
[…]
@@ III.2 Enhancing DMRG
Since the virtual MPS provides direct information on the distribution of entanglement within the simulated quantum circuit section, we can use this structure to construct a grouped MPS by clustering qubits that are highly entangled throughout the evolution. The goal is to group together sites connected by larger bond dimensions, while using the lower ones as the effective bond dimensions of the grouped MPS. At the same time, we should verify that the generated grouped sites actually fit within the available memory. This problem can be solved by mapping the virtual MPS into a graph, where the bond dimensions are represented as edge weights. The task then becomes to find a min-cut multipartitioning of the graph, such that qubits connected by large bond dimensions are grouped within the same partition, while the lower bond dimensions correspond to the edges cut by the partitioning process. To achieve this, we make use of the KaHyPar graph partitioning tool [35] to perform recursive bipartitioning. In the first iteration, we divide the full virtual MPS into two subnetworks, with sites from $1$ to $k_{1}$ and from $k_{1}+1$ to $N$ respectively, and check the following condition for each subnetwork:
@@
with $N_{1}(p)$ being the number of qubits in partition $p$, $\chi_{k}$ the bond dimension that separates the two subnetworks, and $Q_{\mathrm{max}}$ a parameter that limits the size of each grouped site - similar to cluster-TEBD. The number of qubits $N_{1}(p)$ in partition $p$ can be easily estimated as $k_{1}$ for the first partition, and $N-k_{1}+1$ for the second one. If the condition in Eq. (10) does not hold for any of the two subnetwork, the algorithm recurs by applying the same procedure on the corresponding subnetwork. At a generic step $i>1$ of the recursion, the previous condition can be generalized as follows:
@@
with $\chi_{k_{j}}$ being the bond dimension connecting two subnetworks that were separated at a previous recursion step $j~{}<~{}i$. The algorithm stops recurring when Eq. (11) - or Eq. (10) for subnetworks at the extremes of the MPS - holds for all the generated subnetworks. Finally, we use the obtained partitioning of the virtual MPS as our grouping scheme. The sites of the initial state belonging to the same partition $\tau$ are grouped in the corresponding site $\tau$, forming the grouped MPS that will be used in the iteration. Fig. 7 depicts an example of this grouping algorithm applied on the circuit from Fig. 1.
@@ IV.1 Tested quantum circuits
We begin by considering a circuit structure inspired by random circuit sampling [16, 18], where each circuit not only contains random gates but it is also assembled randomly, similar to the design of local random quantum circuits [49, 50, 51, 52, 53, 54]. We refer to this circuit as a “random-structure quantum circuit”. After choosing the number of qubits and layers, a random-structure quantum circuit is generated by applying either a random single-qubit gate or a random two-qubit gate with a uniform probability $p=0.5$, until all layers are completed.
Figure 8: Comparison of cluster-TEBD and TEBD in simulations of Clifford random-structure quantum circuits. (a) Runtime, in logarithmic scale, of standard TEBD and cluster-TEBD simulations of Clifford random-structure quantum circuits with maximum bond dimension $\chi_{\mathrm{max}}=4096$. (b) Runtime, in logarithimc scale, of standard TEBD and cluster-TEBD simulations of $1000$-qubit Clifford random-structure quantum circuits with bond dimension $4096$ and varying number of layers. (c) Runtime, in logarithmic scale, of standard TEBD and cluster-TEBD simulations of $100$-layer Clifford random-structure quantum circuits with bond dimension $4096$ and varying number of qubits. (d) Fidelity of standard TEBD and cluster-TEBD simulations of $100$-layer Clifford random-structure quantum circuits with varying number of qubits.
The initial state is simply a product of |$\ket{0}$s, for all the qubits in the system. We consider two variations of these random-structure quantum circuits: (i) in the first version, single-qubit gates are randomly selected from the set $\left\{H,X,Y,Z\right\}$ and two-qubit gates from the set $\left\{CNOT,C\text{-}Y,C\text{-}Z,SWAP\right\}$. Since all these gates belong to the Clifford group, all the circuits are stabilizer circuits [37]; (ii) in the second version, the sets are extended, adding the non-Clifford gates $\left\{T,P(\frac{3\pi}{4}),\sqrt[4]{X},\sqrt{W}\right\}$ to the set of single-qubit gates, with $\sqrt{W}=\frac{1}{\sqrt{2}}\left(X+Y\right)$, and the gates $\left\{C\text{-}H,C\text{-}S,C\text{-}T,\sqrt{SWAP}\right\}$ to the set of two-qubit gates. This version introduces a $50\%$ ratio of non-stabilizer elements in the circuits, increasing the complexity of the simulations [37, 55]. For a more detailed description of the gates employed in the simulated circuits, see Appendix A.
@@
While random-structure quantum circuits are dense, with layers filled with gates, this implementation of Shor’s algorithm features a significant number of layers with very few gates per layer. Moreover, gates within similar ranges of layers are localized, meaning they operate on subsets of qubits that are in close proximity, leaving the remaining qubits of the system unaffected. Our approach of clustering together qubits entangled by the evolution proves particularly advantageous in this context of sparse layers, as it is sufficient to focus on a localized section of the state. The broader the range of layers in which these localized operations occur, the greater the number of layers that can be simulated simultaneously within a single cluster. Let us stress that this structure is not a specific feature of Shor’s algorithm, but it appears in other known quantum algorithms [58, 31, 59].
We run all our benchmarks on the CINECA supercomputer LEONARDO, where we employ one node for each simulated circuit and for each algorithm [60]. In particular, we employ Booster nodes, each with 32 cores and 512GB of RAM, where all resources are exclusively dedicated to each simulation to guarantee reliable measures for the runtime.
IV.2 Cluster-TEBD vs. standard TEBD
In this section, we compare the performances of cluster-TEBD and standard TEBD algorithms in terms of runtime and fidelity for the two aforementioned variations of random-structure quantum circuits, i.e., with Clifford and non-Clifford gates respectively.
IV.2.1 Clifford random-structure quantum circuits
We start by showing the results for Clifford random-structure quantum circuits of various sizes. More specifically, we select a number of qubits spanning from $40$ up to $1000$, and circuit depths of $40$, $60$, $80$, and $100$ layers. In the simulations, we vary the maximum bond dimension $\chi_{\mathrm{max}}$ of the MPS from $128$ to $4096$, while the parameter $Q_{\mathrm{max}}$ is set at $27$.

### v2

@@ III Entanglement clustering in finite-fidelity simulations
The core idea of this work is to improve the algorithms for simulating quantum circuits described in the previous section using cluster-based protocols. By leveraging the properties of a given circuit’s structure, we can “cluster” qubits based on regions of the circuit with a large number of entangling gates. Previous work has proposed strategies that group gates into clusters [Fatima2021GateCluster, Chen2021CircuitPartitioner, cicero2024simulationquantumcomputersreview]. In our approach, which we refer to as “entanglement clustering”, clusters are formed by grouping gates applied up to a certain layer on a subset of the total system. Gates within a cluster are not contracted immediately; instead, they are combined with the corresponding portion of the initial state into a tensor network, which is contracted exactly once the cluster reaches its maximal size. The termination criterion for each cluster depends on the size of the resulting tensor, accounting both for the arrangement of multiqubit gates in the analyzed circuit portion and the bond dimension of the state at that stage. This physics-informed clustering approach allows TEBD to adaptively divide the circuit into multilayered clusters, employing optimized tensor network contractions [Gray2021HyperExact] to contract each cluster exactly. In DMRG, a similar approach is applied through an adaptive clustering algorithm, which exploits hypergraph partitioning [Schlag2022KaHyPar] to automatically form clusters of qubits, based on the distribution of entangling operations within a specified segment of the circuit of up to $L_{\mathrm{max}}$ gate layers. This qubit grouping is used to generate a grouped MPS, whose sites are variationally optimized in each iteration. Our implementation of the DMRG algorithm is adaptive, as the structure of the grouped MPS dynamically changes at each iteration to match the entangling gate distribution in the local circuit region.
@@
Figure 6: Schematic of the iteration process of the cluster-TEBD on an eight-qubit, seven-layer quantum circuit, with $Q_{\mathrm{max}}=5$. (a) Up to layer $2$, the circuit forms four clusters of entanglement, with two qubits each. At layer $3$, the operation $U_{(6,7),3}$ groups clusters 33 and $4$ in a new cluster 33 with four qubits. As the size of this cluster does not violate the condition on $Q_{\mathrm{max}}$, we move on to the next layer. (b) At layer $4$, operations $U_{(1,2),4}$ and $U_{(5,6),4}$ are performed on pairs of qubits that are already in the same cluster. Consequently, the clusters remain the same and we move on to the next layer. (c) At layer $5$, the operation $U_{(4,5),5}$ groups clusters 22 and 33 into a new cluster 22 with six qubits. As the size of this cluster violates the condition on $Q_{\mathrm{max}}$, layer $4$ is the chosen layer up to which the circuit can be contracted exactly. (d) Schematic of the division of clusters up to layer $4$, contraction and decomposition – with truncation, if necessary. The obtained MPS is used as the initial state of the remaining evolution.
@@ III.1 Cluster-TEBD
The cluster-TEBD algorithm modifies the TEBD algorithm described in Sec. II.3 to allow the exact contraction of multiple gate layers of a quantum circuit. The idea is to divide the network into clusters of entanglement, based on the structure of the circuit to simulate. These clusters are independent from each other, meaning they can be interpreted as intermediate-to-large-sized tensor networks that can be exactly contracted in an efficient way.
@@
The idea of gate clustering has already been successfully applied in the context of exact simulations of quantum circuits [Fatima2021GateCluster, Chen2021CircuitPartitioner, cicero2024simulationquantumcomputersreview], and our purpose is to exploit it in finite-fidelity tensor network algorithms.
@@
In the context of exact tensor network contractions, a similar approach is the “slicing” technique [Gray2021HyperExact, Chen2018classicalsimulationintermediatesizequantum, Villalonga2019hpcsimqc, Huang2020classical, Schutski2020SimpleHeuristics, zhang2019alibabacloudquantumdevelopment, Huang2021Slicing, cicero2024simulationquantumcomputersreview, tang2025tensorqcscalabledistributedquantum, herzog2025jointcuttinghybridschrodingerfeynman, sarode2025circuitpartitioningcircuitexecution], where a network is divided into independent subnetworks by choosing a subset of indices to fix. As a consequence, the contraction of all subnetworks has to be repeated multiple times, for every possible value of the fixed indices. The main difference in our approach is that no indices within the network are fixed, eliminating the need to repeat the contraction of the generated subnetworks. Instead, the circuit partitioning is only based on the entanglement distribution and on the available memory, making sure that, after all subnetworks have been contracted, the resulting tensor sizes are kept manageable for the SVD compression step.
@@
A single iteration of cluster-TEBD is performed as follows. We define an array of length N1N-1, representing bonds between neighboring qubits, with all entries initially set to False. Each entry $i$ in the array corresponds to the bond between qubits $i$ and $i+1$, and its value indicates whether an entangling operation has been applied between them. The algorithm iterates over all gates in the circuit, starting from the layer closest to the initial state, to track the distribution of entangling operations. When encountering an $n$-qubit gate, with $n>1$, acting on qubits $i,\dots,i+n-1$, the corresponding entries $i,\dots,i+n-2$ of the bond array are updated to True. This signifies that an entangling operation has been applied between these qubits. At the end of each layer, the algorithm identifies entanglement clusters by finding consecutive entries in the bond array that are set to True. In the context of the cluster-TEBD algorithm, an entanglement cluster can be interpreted as a portion of the circuit, up to the current layer $L_{\text{current}}$, where qubits frequently interact with each other but never interact with the rest of the system. Since distinct clusters of entanglement can be treated independently, the corresponding circuit portion can be contracted as a separate circuit. The size of the tensor resulting from the contraction of a cluster depends on two quantities:
@@
Knowing that all the physical indices of a quantum circuit have dimension $2$, the size of tensor $\Psi^{[k]}_{L_{\text{current}}}$ that exactly encodes the entanglement cluster $k$ up to layer $L_{\text{current}}$ can be computed as
@@
The quantum circuit can then be contracted exactly up to the last layer $L$ that satisfies both Eqs. (7) and (8). Once layer $L$ has been identified, the algorithm proceeds to exactly contract all entanglement clusters, dividing each portion of the circuit from the initial state up to $L$ into independent tensor networks. The exact contraction of each cluster $k$ exploits the greedy algorithm to find the corresponding contraction sequences [Smith2018opt_einsum, Gray2021HyperExact]. The resulting tensors $\Psi_{L}^{[k]}~\in~\mathbb{C}^{\chi_{q_{1}^{[k]}-1}\times q_{1}^{[k]}\times\dots\times q_{n}^{[k]}\times\chi_{q_{n}^{[k]}}}$ are finally decomposed using successive SVDs – as shown back in Fig. 2.
@@
Figure 7: Dynamical adaptive grouping routine for DMRG on an eight-qubit quantum circuit, setting $L_{\mathrm{max}}=7$. (a) Generation of the virtual MPS from the quantum circuit by counting the number of two-qubit gates per bond and estimating the bond dimensions. (b) Recursive bipartitioning protocol for the virtual MPS, setting $Q_{\mathrm{max}}=4$. The generated partitioning is finally used to group the initial state accordingly, creating the grouped MPS for the DMRG algorithm.
III.2 Dynamical Adaptive Grouping Routine for DMRG
The idea of clustering together qubits based on the distribution of entanglement in a quantum circuit can also be used to improve the DMRG algorithm from Sec. II.3. As opposed to regular, periodically invariant circuit structures, for example, a brick-wall structure with alternating two-qubit gates, generic circuits with irregular structures display an uneven distribution of entanglement throughout the evolution. Depending on the time step, distinct subsets of qubits may experience numerous entangling operations, while other separate subsets may undergo very few or even no operations. Our goal is to develop entanglement-based protocols that enhance the applicability of DMRG for simulating generic quantum circuits.
In Sec. II.3, we introduced the concept of a grouping scheme, designed to group together highly entangled sites in the simulated quantum circuit. This idea of forming a grouped MPS is conceptually similar to creating clusters of entanglement, with one key difference: in the original DMRG algorithm, the grouping was predetermined prior to the algorithm’s execution. While this method works well for circuits with regular structures, it becomes less effective for deep circuits with irregularly arranged entangling gates, where it is not straightforward to predict how the bond dimensions will grow in different portions of the circuit. Therefore, we introduce a clustering approach to adaptively choose the site grouping and the needed bond dimensions, according to the structure of the simulated circuit.
An iteration of the dynamical adaptive grouping routine works as follows. Given a generic $N$-qubit quantum circuit, we count the number of entangling operations $E_{L_{\mathrm{max}}}(b)$ for each bond $b$, up to the chosen number of contracted layers per iteration $L_{\mathrm{max}}$. Then, we create a virtual MPS with $N$ sites, where each bond dimension is defined as
$\tilde{\chi}_{b}:=\min\left\{2^{E_{L_{\mathrm{max}}}(b)}\chi_{b},\,\chi_{\mathrm{max}}^{\mathrm{DMRG}},\,2^{\min\left\{b,N-b\right\}}\right\}.$ (9)
@@ III.2 Dynamical Adaptive Grouping Routine for DMRG
Since the virtual MPS provides direct information on the distribution of entanglement within the simulated quantum circuit section, we can use this structure to construct a grouped MPS by clustering qubits that are highly entangled throughout the evolution. The goal is to group together sites connected by larger bond dimensions, while using the lower ones as the effective bond dimensions of the grouped MPS. At the same time, we should verify that the generated grouped sites actually fit within the available memory. This problem can be solved by mapping the virtual MPS into a graph, where the bond dimensions are represented as edge weights. The task then becomes to find a min-cut multipartitioning of the graph, such that qubits connected by large bond dimensions are grouped within the same partition, while the lower bond dimensions correspond to the edges cut by the partitioning process. To achieve this, we make use of the KaHyPar graph partitioning tool [Schlag2022KaHyPar] to perform recursive bipartitioning. In the first iteration, we divide the full virtual MPS into two subnetworks, with sites from $1$ to $k_{1}$ and from $k_{1}+1$ to $N$, respectively, and check the following condition for each subnetwork:
@@
with $N_{1}(p)$ being the number of qubits in partition $p$, $\chi_{k}$ the bond dimension that separates the two subnetworks, and $Q_{\mathrm{max}}$ a parameter that limits the size of each grouped site – similar to cluster-TEBD. The number of qubits $N_{1}(p)$ in partition $p$ can be easily estimated as $k_{1}$ for the first partition, and $N-k_{1}+1$ for the second one. If the condition in Eq. (10) does not hold for any of the two subnetworks, the algorithm recurs by applying the same procedure on the corresponding subnetwork. At a generic step $i>1$ of the recursion, the previous condition can be generalized as follows:
@@
with $\chi_{k_{j}}$ being the bond dimension connecting two subnetworks that were separated at a previous recursion step j<ij~<~i. The algorithm stops recurring when Eq. (11) – or Eq. (10) for subnetworks at the extremes of the MPS – holds for all the generated subnetworks. Finally, we use the obtained partitioning of the virtual MPS as our grouping scheme. The sites of the initial state belonging to the same partition $\tau$ are grouped in the corresponding site $\tau$, forming the grouped MPS that will be used in the iteration. Fig. 7 depicts an example of the dynamical adaptive grouping routine applied on the circuit from Fig. 1.
@@ IV.1 Tested quantum circuits
We begin by considering a circuit structure inspired by random circuit sampling [Aaronson2017QSComplexity, Bouland2019RCS], where each circuit not only contains random gates but it is also assembled randomly, similar to the design of local random quantum circuits [Fisher2023RSQC, Harrow2009RSQC2, Brandao2014LocalRQC, mitsuhashi2024unitarydesignssymmetriclocal, mitsuhashi2024characterizationrandomnessquantumcircuits, suzuki2024globalrandomnessrandomlocal]. We refer to this circuit as a “random-structured quantum circuit”. After choosing the number of qubits and layers, a random-structured quantum circuit is generated by applying either a random single-qubit gate or a random two-qubit gate with a uniform probability $p=0.5$, until all layers are completed.
Figure 8: Comparison of cluster-TEBD and TEBD in simulations of Clifford random-structured quantum circuits. (a) Runtime, in logarithmic scale, of standard TEBD and cluster-TEBD simulations of Clifford random-structured quantum circuits with maximum bond dimension $\chi_{\mathrm{max}}=4096$. (b) Runtime, in logarithmic scale, of standard TEBD and cluster-TEBD simulations of $1000$-qubit Clifford random-structured quantum circuits with bond dimension $4096$ and varying number of layers. (c) Runtime, in logarithmic scale, of standard TEBD and cluster-TEBD simulations of $100$-layer Clifford random-structured quantum circuits with bond dimension $4096$ and varying number of qubits. (d) Fidelity of standard TEBD and cluster-TEBD simulations of $100$-layer Clifford random-structured quantum circuits with varying number of qubits.
@@
[…]
While random-structured quantum circuits are dense, with layers filled with gates, this implementation of Shor’s algorithm features a significant number of layers with very few gates per layer. Moreover, gates within similar ranges of layers are localized, meaning they operate on subsets of qubits that are in close proximity, leaving the remaining qubits of the system unaffected. Our approach of clustering together qubits entangled by the evolution proves particularly advantageous in this context of sparse layers, as it is sufficient to focus on a localized section of the state. The broader the range of layers in which these localized operations occur, the greater the number of layers that can be simulated simultaneously within a single cluster. Let us stress that this structure is not a specific feature of Shor’s algorithm, but it appears in other known quantum algorithms [Chen2023QFTSmallEntanglement, Niedermeier2024QalgsTN, Arnault2024typologyquantumalgorithms].
We run all our benchmarks on the CINECA supercomputer LEONARDO, where we employ one node for each simulated circuit and for each algorithm [LeonardoCINECA]. In particular, we employ Booster nodes, each with 32 cores and 512GB of RAM, where all resources are exclusively dedicated to each simulation to guarantee reliable measures for the runtime.
IV.2 Cluster-TEBD versus standard TEBD
In this section, we compare the performances of cluster-TEBD and standard TEBD algorithms in terms of runtime and fidelity for the two aforementioned variations of random-structured quantum circuits, i.e., with Clifford and non-Clifford gates, respectively.
IV.2.1 Clifford random-structured quantum circuits
We start by showing the results for Clifford random-structured quantum circuits of various sizes. More specifically, we select a number of qubits spanning from $40$ up to $1000$, and circuit depths of $40$, $60$, $80$, and $100$ layers. In the simulations, we vary the maximum bond dimension $\chi_{\mathrm{max}}$ of the MPS from $128$ to $4096$, while the parameter $Q_{\mathrm{max}}$ is set at $27$.